OpenStudy (anonymous):

divide and simplify. x^2-81/x\[\div\]x-9/x+7

OpenStudy (aroub):

x^2-81 x-9 ------- ÷ ---------- x x+7 = (x-9)(x+9) x+7 ---------- * -------------- x x-9 = (x+9) x+7 ------ * ---------- x 1 = (x+9) (x+7) ------------ x \[x \neq 0 \]